F(x)= x^3-5x^2+144x-720
I got one of the zeroes (x=5) I need the other 2... they should be imaginary numbers.

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

y = (x-5) (x^2+144)
you should able to figure this out
x^2+144=0
x^2=-144
and you are right there is an i in there.

Answer by Alan3354(69443)

F(x)= x^3-5x^2+144x-720
I got one of the zeroes (x=5) I need the other 2
-----------------
Divide by (x-5):
--> x^2 + 144
x = ± 12i